At what potential difference must an X-ray tube operate to produce X-rays with a
minimum wavelength of 2A. °
1
Expert's answer
2019-04-05T10:26:08-0400

From the conservation of energy:

"eU=\frac{hc}{λ}"

"U=\frac{hc}{eλ}=\frac{1239.84 eVnm}{(1e)(0.2 nm)}=6200 V=6.2 kV"